which is an equation for the line that contains the points (-2, 3) and (2, -1)?

Answer by Nate(3500)

(-2, 3) and (2, -1)
the equation for slope is: (y(sub2)-(y(sub1))/((x(sub2)-(x(sub1))
use (-2 (sub1), 3(sub1)) and (2(sub2), -1(sub2))
(y(sub2)-(y(sub1))/((x(sub2)-(x(sub1))
(-1-3)/(2-(-2))
(-4)/4
the slope is -1
the equation is: y - (y-point)=m(x-(x-point)) where m is the slope
y - (y-point)=m(x-(x-point)) original; let us use (-2,3)
y - 3 = -1(x - (-2))
y-3=-1(x+2)
y-3=-x-2
y=-x+1
the equation is y=-x+1